# 22-descent for Bloch–Kato Selmer groups and rational points on hyperelliptic curves I

### v3

@@ 8. $2$-adic Coleman integration
    $y_{n}$ is a lift of $\phi(y)$ modulo $2^{n}$.
@@
    The lift $\phi$ defined by sending $x$ to $x^{2}$ and $y$ to $\lim_{n}y_{n}$ extends to a function on $\{|x|<2^{6}\}$.
Proof.
We have
(19) $Q(x^{2},y_{n+1})=2Q(x^{2},y_{n})-2y_{n}Q(x^{2},y_{n})+Q(x^{2},y_{n})^{2}.$
This implies that $2^{n}$ divides $Q(x^{2},y_{n})$, and hence implies the first part of the lemma. We may write $y_{n}=f_{1}(x)+yf_{2}(x)$ for unique $f_{1},f_{2}\in\mathbb{Z}_{2}[x]$. Then [DV06, Lemma 1] implies that the degree of $f_{1}$ and $f_{2}$ is at most 6(n1)+26(n-1)+2, which implies part (2) of the Lemma. ∎
Lemma 33.
  1. (1)
    If $f$ has odd degree, then the differentials $x^{i}\cdot ydx$, for $0\leq i\leq 2g$, span $H^{1}_{\mathrm{dR}}(X/K).$
  2. (2)
    If $f=x^{5}-x$ then a basis of $H^{1}_{\mathrm{dR}}(X/K)$ is given by $\{x^{i}\cdot ydx:0\leq i\leq 3\}$.
Proof.
The first claim can be seen by changing coordinates to obtain a curve of the form
$y^{2}=g(x),$
using the fact that $H^{1}_{\mathrm{dR}}(X/\mathbb{Q})$ is isomorphic to
$(\mathbb{Q}[x]\cdot dx/y)/\langle(2ix^{i-1}g(x)+g^{\prime}(x))dx/y:i\geq 1\rangle.$
The second claim follows by a similar calculation. ∎
A reduction algorithm for differentials on $X$ with respect to the spanning set $\{x^{i}ydx\}$ above, amounts to an algorithm which, given a differential of the second kind $g(x,y)dx$, returns constants $\lambda_{i}$ in $K$ and a function $h(x,y)\in K(X)$ such that
$g(x,y)dx=\sum_{i}\lambda_{i}x^{i}ydx+dh.$
We have
$d(x^{i}(2y-1)^{3})$ $=(ix^{i-1}(2y-1)^{3}+6x^{i}(2y-1)f^{\prime}(x))dx$
$=2y((3+4f(x))ix^{i-1}+6x^{i}f^{\prime}(x))dx+((4f(x)-1)ix^{i-1}-6x^{i-1}f^{\prime}(x))dx,$
and hence, if $f$ is monic of degree $d$, we get, for any $i\geq 0$,
(20) $x^{d+i-1}ydx=\frac{1}{6d+4i}g_{i}(x)ydx+\frac{1}{6d+8i}(h_{i}(x)dx+d(x^{i}(2y-1)^{3})),$
where $g_{i},h_{i}\in\mathbb{Z}[x]$ have degree less than d+i1d+i-1. An explicit convergence estimate is given in [DV02, Lemma 1].
8.1. The universal unipotent connection on $X$
Given a point $b\in X(K)$, we say that a vector bundle with flat connection $\mathcal{E}_{n}$ on $X$ together with a vector $v\in b^{*}\mathcal{E}_{n}$ is a universal $n$-unipotent connection if for every $n$-unipotent vector bundle with connection $\mathcal{V}$ and vector $v$ in $b^{*}\mathcal{V}$, there exists a unique morphism $\mathcal{E}_{n}\to\mathcal{V}$ sending $e_{n}$ to $v$. The vector bundle with connection $\mathcal{E}_{1}$ is an extension of $\mathcal{O}_{X}$ by $\mathcal{O}_{X}\otimes H^{1}_{\mathrm{dR}}(X/K)^{*}$. The vector bundle $\mathcal{E}_{2}$ is an extension of $\mathcal{E}_{1}$ by
$\mathcal{O}_{X}\otimes\mathrm{Ker}(H^{1}_{\mathrm{dR}}(X/K)\otimes H^{1}_{\mathrm{dR}}(X/K)\stackrel{{\scriptstyle\cup}}{{\longrightarrow}}H^{2}_{\mathrm{dR}}(X/K))^{*}.$
For any $z\in X(\mathbb{F}_{2})$, the connection $\mathcal{E}_{n}$ admits a rigid analytic trivialisation upon restriction to the residue disk ]z[]z[, i.e. the map
$(\mathcal{E}_{n}(]z[)^{\nabla=0}\otimes_{\mathbb{Q}_{2}}\mathcal{O}_{]z[},d)\to(\mathcal{E}_{n}|_{]z[},\nabla)$
is an isomorphism of connections. In particular, for all $\widetilde{z}\in]z[(\mathbb{Q}_{2})$ the evaluation map
$\mathrm{ev}_{\widetilde{z}}:\mathcal{E}_{n}(]z[)^{\nabla=0}\to\widetilde{z}^{*}\mathcal{E}_{n}$
is an isomorphism. For $z_{1},z_{2}\in]z[(\mathbb{Q}_{2})$ we define $I_{z_{1},z_{2}}$ to be the parallel transport isomorphism
$\mathrm{ev}_{z_{2}}\circ\mathrm{ev}_{z_{1}}^{-1}:z_{1}^{*}\mathcal{E}_{n}\to z_{2}^{*}\mathcal{E}_{n}.$
8.1.1. Frobenius structure
To compute the Frobenius structure on $\mathcal{E}_{n}$ amounts to computing an isomorphism
$F:\mathcal{E}_{n}\simeq\phi^{*}\mathcal{E}_{n}$
of isocrystals, which amounts to computing a matrix $G$ satisfying
$\phi^{*}\Lambda\cdot G-G\cdot L-dG=0$
Given a lift $\phi$ of Frobenius on ]U[]U[, for all $x\in U(\mathbb{F}_{2})$ there is a unique point $\iota(x)\in]U[(\mathbb{Q}_{2})$ fixed by $\phi$, by Hensel’s lemma. Given an isocrystal $\mathcal{E}$ with Frobenius structure $F$, and a point $\overline{z}\in U(\mathbb{F}_{2})$ we hence obtain an automorphism $z_{0}^{*}(F)$ of $z_{0}^{*}\mathcal{E}$, where $z_{0}:=\iota(\overline{z})$. For general $z$ in ]U[]U[, we define $z^{*}F$ to be the automorphism $I_{z_{0},z}\circ z_{0}^{*}F\circ I_{z,z_{0}}$.
By definition we have the following characterisation of the Frobenius action on $\mathcal{E}_{n}$.
Lemma 34.
Let $b\in X(\mathbb{Q}_{2})$ be a point not reducing to the point at infinity in $X(\mathbb{F}_{2})$. Then the Frobenius structure on $\mathcal{E}_{n}$ associated to the basepoint $b$ is the unique Frobenius structure $\Phi_{b}$ such that the map
$1:\mathbb{Q}_{p}\to b^{*}\mathcal{E}_{n}$
is $\phi$-equivariant.
Given points $z_{1},z_{2}$ congruent modulo $p$, we define $J_{z_{1},z_{2}}$ to the parallel transport isomorphism
$z_{1}^{*}\mathcal{E}\stackrel{{\scriptstyle\simeq}}{{\longrightarrow}}z_{2}^{*}\mathcal{E}$
defined by the connection on $\mathcal{E}$. Explicitly, let $t$ be a parameter at $z_{1}$ and let $\mathbb{Q}_{p}[\![t]\!]$ be the formal completion of the local ring at $z_{1}$. Then take $J_{z_{1}}$ to be the unique matrix in $1+t\mathrm{Mat}_{n}(\mathbb{Q}_{p}[\![t]\!])$ satisfying
$\Lambda J_{z_{1}}=dJ_{z_{1}}.$
Then the Frobenius structure on $\Lambda$ ensures that $J_{z_{1}}$ converges on the open disk of radius $1$ and we define $J_{z_{1},z_{2}}$ to be the evaluation at $z_{2}$.
Given $G$, the Frobenius action $\phi$ at a point $z$ is defined to be the inverse of the endomorphism
(21) $\Phi(z)=J_{\phi(z),z}\circ G(z).$
Hence, by definition, the condition in the Lemma amounts to the identity
(22) $J_{\phi(b),b}\circ G_{b}(b)(1)=1.$
8.1.2. Hodge filtration
The connection $\mathcal{E}_{2}$ has the structure of a filtered connection, i.e. an exhaustive and separating decreasing filtration $F^{i}\mathcal{E}$ satisfying the Griffiths transversality condition
$\nabla(F^{i}\mathcal{E})\subset F^{i-1}\mathcal{E}\otimes\Omega_{X}.$
To compute the Hodge filtration on $\mathcal{E}_{2}(X)$, we could apply the algorithm from [BD21], which describes how to compute the Hodge filtration for a hyperelliptic curve of the form $y^{2}=f(x)$. This gives a description of the sub-bundles $F^{i}\mathcal{E}_{2}(X)$ in terms of a local trivialisation of $\mathcal{E}_{2}(X)$ using a basis of $H^{1}_{\mathrm{dR}}(X)$ of the form $x^{i}dx/y$. As we use a different basis for Frobenius calculations, it is simpler to redo the whole calculation.
Theorem 5 (Hadian).
The Hodge filtration $F^{i}\mathcal{E}_{n}$ is uniquely determined by the following conditions.
  1. (1)
    The sequence
    $(H^{1}_{\mathrm{dR}}(X/K)^{\otimes n})^{*}\otimes\mathcal{O}_{X}\to\mathcal{E}_{n}(X)\to\mathcal{E}_{n-1}(X)\to 0$
    is strictly compatible with the Hodge filtrations on $\mathcal{E}_{n}(X),\mathcal{E}_{n-1}(X)$ and $H^{1}_{\mathrm{dR}}(X/K)$.
  2. (2)
    $e_{n}\in b^{*}F^{0}\mathcal{E}_{n}(X)$.
We deduce that to compute the Hodge filtration, it is essentially enough to compute the extension of $\mathcal{E}_{2}(X)$ to a connection on $X$. Putting the last two subsubsections together, we have enriched the flat connection $\mathcal{E}_{n}$ with the structure of a filtered $F$-isocrystal over the smooth model $X/\mathbb{Z}_{2}$.
8.1.3. Extending to infinity
For our final calculations it will be important to compute the generalised height, and hence the matrix of Frobenius, at the residue disk at infinity. For this, we use overconvergence of our lift of Frobenius. Namely, if our Frobenius lift is defined on a strict neighbourhood $U$ of ]Y[]Y[ for an open subscheme $Y\subset X_{\mathbb{F}_{p}}$ (assume for simplicity that all points of XYX-Y are defined over $\mathbb{F}_{p}$), we may apply (21) to deduce that for any $z_{1},z_{2}\in]x[(\overline{\mathbb{Q}}_{p})$, we have
(23) $\Phi(z_{1})=J_{\phi(z_{2}),z_{1}}\circ G(z_{2})\circ J_{z_{1},z_{2}}.$
If $z\in(X-Y)(\mathbb{F}_{p})$, then for any $z_{1}\in]z[(\mathbb{Q}_{p})$ and $z_{2}\in]z[\cap U(\overline{\mathbb{Q}}_{p})$, we can compute $G(z_{2})$ and use (23) to obtain a formula for $\Phi(z_{1})$. See [BT20] for a related construction in the context of abelian Coleman integrals.
Note that typically $U\cap]z[$ will be an annulus with no $\mathbb{Q}_{p}$-points, and hence we will have to work over a ramified extension of $\mathbb{Q}_{p}$.
8.2. Computing the generalised height
We are now ready to define our function from $\mathbb{Q}_{p}$-points to filtered $\phi$-modules as follows. We have a short exact sequence of filtered $F$-isocrystals
$0\to(H^{1}_{\mathrm{dR}}(X/\mathbb{Q}_{p})^{*\otimes 2}/\mathbb{Q}_{p}(1))\otimes_{\mathbb{Q}_{p}}\mathcal{O}_{X}\to\mathcal{E}_{2}\to\mathcal{E}_{1}\to 0.$
Here the connection on the left hand side is constant, and the Frobenius and Hodge filtration is induced from that on $H^{1}_{\mathrm{dR}}(X/\mathbb{Q}_{p})$. We define
$\mathcal{E}:=\mathcal{E}_{2}/\mathrm{Sym}^{2}H^{1}_{\mathrm{dR}}(X/\mathbb{Q}_{p})^{*}\otimes\mathcal{O}_{X}.$
For each $\in X(\mathbb{Q}_{p}), xx^{*}\mathcal{E}$ then has the structure of a mixed extension of filtered $\phi$-modules with graded pieces $\mathbb{Q}_{p},H^{1}_{\mathrm{dR}}(X/\mathbb{Q}_{p})^{*}$ and $\wedge^{2}H^{1}_{\mathrm{dR}}(X/\mathbb{Q}_{p})^{*}/\mathbb{Q}_{p}(1)$.
We define the $p$-adic local generalised height of $(b,z)$ (with respect to a splitting of the Hodge filtration of $H^{1}_{\mathrm{dR}}(X/\mathbb{Q}_{p})$) to be the local generalised height of $x^{*}\mathcal{E}$ in the sense of [BD21]. Rather than recall the general definition, we recall an explicit formula.
Lemma 35 ([BD21], Lemma 3.9).
Let $M$ be a filtered $\phi$-module with graded pieces $\mathbb{Q}_{p},V_{\mathrm{dR}},W_{\mathrm{dR}}$. Let
$\rho_{\phi}:\mathbb{Q}_{p}\oplus V_{\mathrm{dR}}\oplus W_{\mathrm{dR}}\simeq M$
and
$\rho_{\mathrm{fil}}:\mathbb{Q}_{p}\oplus V_{\mathrm{dR}}\oplus W_{\mathrm{dR}}\simeq M$
be unipotent isomorphisms respecting the Frobenius actions and Hodge filtrations respectively. Suppose
$\rho_{\mathrm{fil}}^{-1}\circ\rho_{\phi}=\left(\begin{array}[]{ccc}1&0&0\\ \alpha&1&0\\ \gamma&\beta&1\end{array}\right).$
Then, for any splitting $s:V_{\mathrm{dR}}/F^{0}\to V_{\mathrm{dR}}$ of the Hodge filtration, with associated local generalised height $\widetilde{h}_{p}$, we have
$\widetilde{h}_{p}(M)=\gamma-\beta(\alpha-s\circ\pi(\alpha)).$
Finally, we recall the following explicit formula for the Frobenius equivariant splitting $\rho_{\phi}$ in terms of the Frobenius matrix $\Phi(z)^{-1}$ defined earlier.
Lemma 36.
Let $M$ be a mixed extension of filtered $\phi$-modules with graded pieces $\mathbb{Q}_{p},V_{\mathrm{dR}},W_{\mathrm{dR}}$, where $\mathbb{Q}_{p},V_{\mathrm{dR}}$ and $W_{\mathrm{dR}}$ have pairwise distinct Frobenius eigenvalues. Suppose that, with respect to some vector space splitting
$\rho_{0}:\mathbb{Q}_{p}\oplus V_{\mathrm{dR}}\oplus W_{\mathrm{dR}}\simeq M_{\mathrm{dR}},$
the action of Frobenius on $M$ is given by a block matrix
$G_{M}:=\rho_{0}^{-1}\circ\phi_{M}\circ\rho_{0}=\left(\begin{array}[]{ccc}1&0&0\\ \boldsymbol{f}&F_{2}&0\\ \boldsymbol{h}&\boldsymbol{g}&F_{2}\\ \end{array}\right).$
Then the unique Frobenius equivariant splitting
$\rho_{\phi}:\mathbb{Q}_{p}\oplus V_{\mathrm{dR}}\oplus W_{\mathrm{dR}}$
is given by
$\rho_{\phi}=\rho_{0}\circ\left(\begin{array}[]{ccc}1&0&0\\ \alpha&1&0\\ \gamma&\beta&1\end{array}\right),$
where
$\alpha$ $=(1-F_{1})^{-1}\boldsymbol{f}$
$\beta$ $=(F_{1}-F_{2})^{-1}\boldsymbol{g}$
$\gamma$ $=(1-F_{2})^{-1}(h+g\alpha)$
and $(F_{1}-F_{2})$ is viewed as an endomorphism of $\mathrm{Hom}(V_{\mathrm{dR}},W_{\mathrm{dR}})$ in the obvious way.
@@ 9.1. The BMSST curve
For $v=139$ and $449$, a regular semistable model has special fibre an elliptic curve intersecting itself at one point. Hence all $\mathbb{Q}_{v}$-points lie on a common irreducible component, and hence the local generalised height at $v$ of $E(b,z)$ is zero for all $b,z\in X(\mathbb{Q}_{v})$.
By the description of local generalised heights away from $2$, for each element $D$ of
$\mathrm{Ker}(\mathrm{Div}^{0}(X(\mathbb{Q}))\to\mathrm{Sym}^{2}J(\mathbb{Q})\otimes\mathbb{Q}),$
the generalised height of $D$ defines an element of $\mathrm{Ker}(\wedge^{2}H^{1}_{\mathrm{dR}}(X/\mathbb{Q}_{2})\to H^{2}_{\mathrm{dR}}(X/\mathbb{Q}_{2}))^{*}/F^{0}$ which is in the image of $H^{1}_{f}(\mathbb{Q},\wedge^{2}V/\mathbb{Q}_{p}(1))$. Since $\mathrm{Div}^{0}(S)\otimes\mathbb{Q}$ surjects onto $\mathrm{Sym}^{2}J(\mathbb{Q})\otimes\mathbb{Q}$, the rank of $\mathrm{Ker}(\mathrm{Div}^{0}(S)\to\mathrm{Sym}^{2}J(\mathbb{Q})\otimes\mathbb{Q})$ is two. Using the algorithm for computing the local generalised height at $2$ described above, we can verify that the rank of the image of $\mathrm{Ker}(\mathrm{Div}^{0}(S)\to\mathrm{Sym}^{2}J(\mathbb{Q})\otimes\mathbb{Q})$ in $\mathrm{Ker}(\wedge^{2}H^{1}_{\mathrm{dR}}(X/\mathbb{Q}_{2})\to H^{2}_{\mathrm{dR}}(X/\mathbb{Q}_{2}))^{*}/F^{0}$ under the local generalised height is equal to $2$. Hence we can solve for the equations for $X(\mathbb{Q}_{2})_{2}$ described in Lemma 31. In concrete terms, let $t:X(\mathbb{Q})\to\mathbb{Q}_{2}$ send a point $z$ to the $P_{1}-\{\infty\}$ coefficient of $z-\{\infty\}$ in $J(\mathbb{Q})$ with respect to the basis $P_{1}-\{\infty\},P_{2}-\{\infty\},P_{3}-\{\infty\}$. For $i=1,2$, let
$\log_{J,i}:J(\mathbb{Q}_{2})\to\mathbb{Q}_{2}$
be the map sending $P$ to $\int_{P}\omega_{i}$ for any $\omega_{1},\omega_{2}$ of $H^{0}(X_{\mathbb{Q}_{2}},\Omega_{X|\mathbb{Q}_{2}})$. Let
$\Psi:X(\mathbb{Q}_{2})\to\mathbb{Q}_{2}[T]^{\oplus 3}$
be the map
$z\mapsto(\log_{J,1}(z-\{\infty\}),\log_{J,2}(z-\{\infty\}),T)$
and let $\psi:X(\mathbb{Q})\to\mathbb{Q}_{2}^{3}$ be the map $(\log_{J,1}(z-\{\infty\}),\log_{J,2}(z-\{\infty\}),t(z))$. Let $\mathrm{Sym}^{2}:\mathbb{Q}_{2}^{\oplus 3}\to\mathbb{Q}_{2}^{\oplus 6}$ be the natural map sending an element of $\mathbb{Q}_{p}^{\oplus 3}$ to its symmetric square (with respective to any basis of $\mathrm{Sym}^{2}(\mathbb{Q}_{p}^{\oplus 3})$. Choosing a basis of $\mathrm{Ker}(\wedge^{2}H^{1}_{\mathrm{dR}}(X/\mathbb{Q}_{2})\to H^{2}_{\mathrm{dR}}(X/\mathbb{Q}_{2}))^{*}/F^{0}$, view $\widetilde{h}_{2}$ as a map $X(\mathbb{Q}_{2})\to\mathbb{Q}_{2}^{4}.$ For $z\in X(\mathbb{Q})$, we define $\widetilde{\psi}(z):=\mathrm{Sym}^{2}(\psi(z))\oplus\widetilde{h}_{2}(z),$ and similarly define $\widetilde{\Psi}(z)\in\mathbb{Q}_{2}[T]^{10}$ for $z\in X(\mathbb{Q}_{2})$.
For $z\in X(\mathbb{Q}_{2})$ let $M(z)(T)\in\mathrm{Mat}_{10,9}(\mathbb{Q}_{2}[T])$ denote the matrix $(\widetilde{\psi}(P_{1}),\ldots,\widetilde{\psi}(P_{8}),\widetilde{\Psi}(z))$. Let $F_{1}$ and $F_{2}$ denote the determinants of the submatrices obtained by deleting the $9$th and $10$th rows. If $z\in X(\mathbb{Q})$, then setting $T=t(z)$ we see that the rank of $M(z)(T)$ is $8$, and hence $F_{1}$ and $F_{2}$ have a common zero, i.e. the Coleman function obtained by taking the resultant $\mathrm{Res}(F_{1},F_{2})$ vanishes at $z$.
Since the codimension of $\mathrm{Sel}(U_{2})$ in $H^{1}_{f}(\mathbb{Q}_{2},U_{2})$ is one, in general one would not expect that $X(\mathbb{Q}_{2})_{2}$ is equal to $X(\mathbb{Q})$, and indeed we found that $X(\mathbb{Q}_{2})$ contained extra points. Specifically, we found that
@@
Another interesting phenomenon which arises is that of multiple zeroes of Chabauty–Coleman functions. In the classical Chabauty–Coleman method where one considers the zeroes of a Coleman integral $\int\omega$, since one computes a $p$-adic approximation one cannot computationally distinguish between multiple zeroes of a function and distinct simple zeroes which are congruent beyond the level of $p$-adic precision. Hence as well as applying the Mordell–Weil sieve, one either needs all rational points to constitute simple zeroes of the Chabauty–Coleman function, or to be able to prove that they vanish to the order observed numerically. One example where such multiple zeroes can be proved is when the curve in question is a covering of a curve satisfying the Chabauty–Coleman bound, and the cover is ramified at a rational point.
In our situation, all the rational points are simple zeroes of the function $\mathrm{Res}(F_{1},F_{2})$, except for the point at infinity, which can be proved to have a zero of order 4. Namely, if we write the polynomial $F_{i}(T)$ as $T^{2}+a_{i}(z)\cdot T+b_{i}(z)$, where $a_{i}(z)$ and $b_{i}(z)$ are Coleman functions on $X$, then by construction $a_{i}(z)$ vanishes at infinity, and $b_{i}(z)$ vanishes to order $2$ at infinity, since it vanishes at infinity and is even with respect to the hyperelliptic involution. This implies that the resultant of $F_{1}$ and $F_{2}$ will have a zero of order at least 4 at infinity.
@@
(and their Galois conjugates and Weierstrass involutions). Labelling these points $Q_{i}$, and the known rational points from (25) $P_{i}$, we compute that
$h(Q_{1})-3h(Q_{2})+2h(Q_{3})-6h(Q_{5})-\frac{175}{4}h(P_{1})-3h(P_{2})+h(P_{3})$
$-\frac{23}{28}h(P_{4})+2h(P_{5})+\frac{16}{7}h(P_{8}),$
$\frac{81}{2}h(P_{1})+3h(P_{2})-9h(P_{3})+\frac{9}{14}h(P_{4})-3h(P_{5})-\frac{18}{7}h(P_{6})+h(P_{7})$
and $54h(P_{1})-6h(P_{2})-6h(P_{3})+3h(P_{5})+2h(P_{6})+h(P_{8})$ span a two dimensional subspace of the four dimensional vector space
$\wedge^{2}H^{1}_{\mathrm{dR}}(X/\mathbb{Q}_{2})^{*}/(H^{2}_{\mathrm{dR}}(X/\mathbb{Q}_{2})^{*}+F^{0})$ (modulo a large power of $2$). This is a consequence of our descent calculation, since their divisors vanish in $\mathrm{Sym}^{2}\mathrm{Jac}(X)(\mathbb{Q}(\sqrt{-3}))^{\operatorname{Gal}(\mathbb{Q}(\sqrt{-3})|\mathbb{Q})}$ and hence their generalised heights are in the image of the two dimensional vector space $H^{1}_{f}(\mathbb{Q},V_{2}J/\mathbb{Q}_{2}(1))$.

### v4

@@ 6.4. Generalised heights
    $h_{2}$ is a nonzero Coleman function. In particular, when restricted to each residue disc it is given by a rigid analytic function (i.e. a convergent power series).
@@
    $h_{2}$ is a quadratic function on $X(\mathbb{Q})$, in the sense that if $x_{i}\in X(\mathbb{Q})$ and $\lambda_{i}\in\mathbb{Q}_{2}$ satisfy
    $\sum\lambda_{i}([x_{i}]-[\infty])^{2}=0$
    in $\mathrm{Sym}^{2}J(\mathbb{Q})\otimes\mathbb{Q}$ then
    $\sum\lambda_{i}h_{2}(x_{i})=0.$
    Here $J$ denotes the Jacobian of $X$.
Now suppose that the rank of the $J$ is $r$, $X(\mathbb{Q}_{2})_{1}$ is infinite, and one has found points $x_{1},\ldots,x_{m}\in X(\mathbb{Q})$, where $m=\frac{r(r+1)}{2}$, such that the dimension of $(x_{1}-\infty,\langle h_{2}(x_{1})),\ldots,(x_{m}-\infty,h_{2}(x_{m}))\rangle$ span $\mathrm{Sym}^{2}J(\mathbb{Q})\otimes\mathbb{Q}_{2}$ under the map
$P\mapsto P^{2}.$
Let $\Psi$ denote the map
$X(\mathbb{Q}_{2})\to\mathrm{Sym}^{2}H^{0}(X_{\mathbb{Q}_{2}},\Omega)^{*}\oplus W.$
where the first component sends $z$ to the functional
$\omega_{0}\omega_{1}\mapsto(\int^{z}_{\infty}\omega_{0})(\int^{z}_{\infty}\omega_{1}),$
and the second component is $W$. Let $W_{0}\subset W$ denote the vector space spanned by $\Psi(x_{1}),\ldots,\Psi(x_{m})$. Then $X(\mathbb{Q}_{2})_{2}$ is equal to the preimage of $W_{0}$ in $X(\mathbb{Q}_{2})$.
Remark 1.
In practice, one constructs a generalised height by first constructin a local height (or pre-height)
$\widetilde{h}_{2}:X(\mathbb{Q}_{2})\to H^{1}_{f}(\mathbb{Q}_{2},\wedge^{2}V/\mathbb{Q}_{2}(1))$
with the property that $\widetilde{h}_{2}$ is quadratic on $X(\mathbb{Q})$ (in the sense above) modulo $\mathrm{loc}_{2}H^{1}_{f}(\mathbb{Q},\wedge^{2}V/\mathbb{Q}_{2}(1))$. Hence, to compute $X(\mathbb{Q}_{2})_{2}$ in the above set-up, one needs $\binom{r(r+1)}{2}+\dim\mathrm{loc}_{2}H^{1}_{f}(\mathbb{Q}_{2},\wedge^{2}V/\mathbb{Q}_{2}(1))$ points in $X(\mathbb{Q})$ which are suitably independent.
@@ 7.1. The BMSST curve
For $v=139$ and $449$, a regular semistable model has special fibre an elliptic curve intersecting itself at one point. Hence all $\mathbb{Q}_{v}$-points lie on a common irreducible component, and we are in the setting above.
Details of the computation of $X(\mathbb{Q}_{2})_{2}$ may be found in [Dog26]. Since the codimension of $\mathrm{Sel}(U_{2})$ in $H^{1}_{f}(\mathbb{Q}_{2},U_{2})$ is one, in general one would not expect that $X(\mathbb{Q}_{2})_{2}$ is equal to $X(\mathbb{Q})$, and indeed we found that $X(\mathbb{Q}_{2})$ contains extra points. Specifically,
@@
(and their Galois conjugates and Weierstrass involutions). Using these points, and known rational points, we can find enough divisors to observe the quadratic properties of $h_{2}$.